# Dynamical Mean-Field Theory of Complex Systems on Sparse Directed Networks

### v1

@@ Introduction.
Complex dynamical systems are modeled by $N$ degrees of freedom $x_{i}(t)$ ($i=1,\dots,N$) that evolve in time according to the coupled differential equations
$\dot{x}_{i}(t)=-f(x_{i})+\sum_{j=1}^{N}A_{ij}g(x_{i},x_{j})+\xi_{i}(t),$ (1)
where $\{A_{ij}\}_{i,j=1,\dots,N}$ defines the interaction network and $\xi_{i}(t)$ is an uncorrelated Gaussian noise, with mean zero and variance $\sigma^{2}$. The variable $\xi_{i}(t)$ models external random perturbations, the function $f(x)$ governs the dynamics in the absence of interactions, and the kernel $g(x,x^{\prime})$ shapes the pairwise couplings. Equation (1) models the nonequilibrium dynamics of neural networks Sompolinsky et al. (1988); Kadmon and Sompolinsky (2015); Martí et al. (2018); Crisanti and Sompolinsky (2018); Martorell et al. (2024) and ecosystems Opper and Diederich (1992); Bunin (2017); Altieri et al. (2021), epidemic spreading Mata and Ferreira (2013); Mieghem (2012); Goltsev et al. (2012); Pastor-Satorras et al. (2015), synchronization phenomena Stiller and Radons (1998); Rodrigues et al. (2016); Prüser et al. (2024), opinion dynamics Baumann et al. (2020, 2021), and multivariate Ornstein-Uhlenbeck processes Godrèche and Luck (2018); Crisanti and Paoluzzi (2023); Ferraro et al. (2024). Table 1 specifies $f(x)$ and $g(x,x^{\prime})$ for paradigmatic models of complex behavior.
@@
Ornstein-Uhlenbeck (OU) process Godrèche and Luck (2018) $x$ $x^{\prime}$
@@
Table 1: Explicit form of $f(x)$ and $g(x,x^{\prime})$ in complex systems modeled by Eq. (1).
@@
The foremost problem in the study of complex systems is how to reduce the dynamics of many interacting elements to the dynamics of a few variables Thibeault et al. (2024). Dynamical mean-field theory (DMFT) Martin et al. (1973); Coolen (2001); Hertz et al. (2016) is a powerful method to tackle this problem in the limit $N\rightarrow\infty$, yielding a solution in terms of the path-probability for the effective dynamics of a single degree of freedom. The application of DMFT to models described by Eq. (1) has been attracting an enormous interest Kadmon and Sompolinsky (2015); Martí et al. (2018); Crisanti and Sompolinsky (2018); Roy et al. (2019); Martorell et al. (2024); Bunin (2017); Galla (2018); Schuecker et al. (2018); Altieri et al. (2021); Prüser et al. (2024); Crisanti and Paoluzzi (2023); Poley et al. (2023); Ferraro et al. (2024); Aguirre-López (2024); Park et al. (2024); Pham and Kaneko (2024), especially in the context of neural networks and ecosystems. Phase diagrams of these models reveal a rich phenomenology, including different types of phase transitions Bunin (2017); Martorell et al. (2024), chaotic behavior Sompolinsky et al. (1988); Martí et al. (2018), and coexistence of multiple attractors Bunin (2017); Altieri et al. (2021); Ros et al. (2023). Despite this substantial theoretical progress, DMFT is currently limited to fully-connected networks, where each dynamical variable is coupled to all others by means of Gaussian interaction strengths. Nevertheless, the interactions in real-world complex systems are known to be sparse and heterogeneous Newman (2010); Guimarães (2020). Sparseness indicates that each element of the system interacts on average with a finite number of others, while heterogeneity refers to fluctuations in the local topology of the interaction network. How to integrate these features in the formalism of DMFT for systems modeled by Eq. (1) remains an unresolved challenge, and even basic questions, such as characterizing the phase diagram of sparse complex systems, are still out of reach.
@@
On the other hand, the dynamics of complex systems on sparse networks has been extensively studied in the case of Ising spins Hatchett et al. (2004); Neri and Bollé (2009); Mimura and Coolen (2009); Roudi and Hertz (2011); Aurell and Mahmoudi (2012); Del Ferraro and Aurell (2015); Aurell et al. (2017); Vázquez et al. (2017); Behrens et al. (2023, 2024). In this context, both the cavity method Neri and Bollé (2009) and DMFT Hatchett et al. (2004); Mimura and Coolen (2009) provide an analytic solution in terms of the path-probability for the effective dynamics of a single unit. However, the presence of symmetric interactions induce a temporal feedback, which, in the particular case of sparse systems, leads to an exponential growth of the path-probability dimension Hatchett et al. (2004); Mimura and Coolen (2009); Behrens et al. (2023), rendering numerical computations unfeasible. When the interactions are directed or unidirectional, this issue does not arise, allowing the path-probability equation to be efficiently solved Derrida et al. (1987); Hatchett et al. (2004); Neri and Bollé (2009).
@@
Inspired by these results for Ising spins, in this letter we solve the dynamics of models governed by Eq. (1) on sparse directed networks with an heterogeneous topology. The solution represents a foundational step in the study of sparse complex systems, as it ultimately incorporates the role of the network structure in the formalism of DMFT. Besides that, directed networks are interesting from a practical viewpoint, because they model the non-reciprocal interactions in real-world complex systems Asllani et al. (2018), including the synaptic connections in the human cortex Peng et al. (2024) and the trophic interactions in food-webs Dunne et al. (2002); Bascompte (2009); Guimarães (2020).
@@
By generalizing DMFT to sparse systems, we obtain an exact equation for the path-probability describing the effective dynamics of a single variable for $N\rightarrow\infty$. We solve this equation for different models in table 1 by using the population dynamics algorithm Mézard and Parisi (2001, 2003), which is a fundamental tool in the study of sparse disordered systems Kühn (2008); Metz et al. (2019); Abou-Chacra et al. (1973); Mézard and Montanari (2009). The excellent agreement between our theoretical results and numerical simulations of finite systems confirms that our solution is valid for several models.
@@
As an application, we determine the phase diagram of the neural network model by Sompolinsky et al Sompolinsky et al. (1988) in the sparse regime. We show that the phase diagram displays a trivial fixed-point phase, a nontrivial fixed-point phase, and a chaotic region. Our results for the transition lines, based on the dynamics of macroscopic observables, are consistent with the universal critical lines derived from random matrix theory Neri and Metz (2016, 2020); Metz and Neri (2021). Remarkably, we provide evidence that the transition from the nontrivial fixed-point phase to the chaotic regime coincides with the vanishing of the gap between the leading eigenvalue and the bulk of the network spectrum.
@@
Let $A_{ij}=C_{ij}J_{ij}$ be the elements of the $N\times N$ interaction matrix $\boldsymbol{A}$. The binary variables $C_{ij}\in\{0,1\}$ determine the network topology, while $J_{ij}\in\mathbb{R}$ controls the pairwise interaction strengths. If $C_{ij}=1$, there is a directed edge $j\rightarrow i$ pointing from node $j$ to $i$, while $C_{ij}=0$ otherwise. We also set $C_{ii}=0$. The indegree $K_{i}=\sum_{j=1}^{N}C_{ij}$ and the outdegree $L_{i}=\sum_{j=1}^{N}C_{ji}$ count the number of links entering and leaving node $i$ Newman (2010), respectively. The random variables $\{C_{ij}\}_{i\neq j}$ follow the distribution
@@
where $\mathcal{N}$ is the normalization constant. The degrees $\{K_{i},L_{i}\}_{i=1,\dots,N}$ are independent and identically distributed random variables drawn from $p_{k,\ell}=p_{{\rm in},k}\,p_{{\rm out},\ell}$, where $p_{{\rm in},k}$ and $p_{{\rm out},\ell}$ are the indegree and the outdegree distribution Newman (2010); Metz et al. (2019), respectively. The network parameter $c$ is the average degree
@@
Equation (2) defines a network ensemble where directed links are randomly placed between pairs of nodes with probability $c/N$, subject to the prescribed degree sequences generated from $p_{k,\ell}$. In the limit $N\rightarrow\infty$, the network samples generated from Eq. (2) are equivalent to those produced by the configuration model Newman et al. (2001); Fosdick et al. (2018). The coupling strengths $\{J_{ij}\}_{i,j=1,\dots,N}$ are independent and identically distributed random variables drawn from a distribution $p_{J}$ with mean $\mu_{J}$ and variance $\sigma_{J}^{2}$. The distributions $p_{k,\ell}$ and $p_{J}$ fully specify the network ensemble, allowing for a systematic investigation of how network heterogeneities impact the dynamics of complex systems.
@@
of the probability density $\mathcal{P}[\boldsymbol{x}]$ of observing a dynamical path of states $\boldsymbol{x}(t)=(x_{1}(t),\dots,x_{N}(t))$ in a fixed time interval. The correlation functions of $\{x_{i}(t)\}_{i=1,\dots,N}$ follow from the derivatives of $\mathcal{Z}\left[\boldsymbol{\psi}\right]$ with respect to the external sources $\boldsymbol{\psi}(t)=(\psi_{1}(t),\dots,\psi_{N}(t))$. In particular, the $n$th-moment of the local variable $x_{i}(t)$,
@@
Clearly, the generating functional fulfills $\mathcal{Z}\left[0\right]=1$.
In the supplemental material Metz , we calculate the average of $\mathcal{Z}\left[\boldsymbol{\psi}\right]$ over the network ensemble defined by Eq. (2) for finite $c$, recasting the problem in terms of the solution of a saddle-point integral. More importantly, we give a clear physical interpretation of the order-parameters, simplifying the saddle-point equations and obtaining a feasible solution for the dynamics in the sparse regime. In the limit $N\rightarrow\infty$, the microscopic dynamical variables decouple, and the path-probability $\mathcal{P}[x]$ for the effective dynamics of a single variable $x(t)$ is determined from
@@
×$\times\Bigg{\langle}\delta_{F}\Bigg{[}\dot{x}(t)+f(x(t))-\sum\limits_{j=1}^{k}J_{j}\,g(x(t),x_{j}(t))-\xi(t)\Bigg{]}\Bigg{\rangle}_{\xi},$ (7)
where $\delta_{F}$ is the functional Dirac-$\delta$, and $\langle\dots\rangle_{\xi}$ represents the average over the uncorrelated Gaussian noise $\xi(t)$. The macroscopic observables are given by $m(t)=\langle x(t)\rangle_{*}$ and $q(t)=\langle x^{2}(t)\rangle_{*}$, where
@@
is the average over the effective dynamics governed by $\mathcal{P}[x]$. The self-consistent Eq. (7) is the exact solution of a broad class of models (see table 1) on sparse directed networks with arbitrary distributions $p_{J}$ and $p_{k,\ell}=p_{{\rm in},k}p_{{\rm out},\ell}$. We solve the more general case of networks with correlated indegrees and outdegrees in Metz . The solution of Eq. (7) determines the time-evolution of the full probability distribution of the microscopic variables in the limit $N\rightarrow\infty$.
Figure 1: Dynamics of the mean $m(t)$ and the standard deviation $\sqrt{q(t)-m^{2}(t)}$ for the SIS model, LV model, and NN model (see table 1) without external noise ($\sigma=0$). Results are shown for directed random networks with average degree $c=5$ and two indegree distributions (Eq. (9)): Poisson (left column) and geometric (right column). The distribution $p_{J}$ of the coupling strengths has mean $\mu_{J}$ and standard deviation $\sigma_{J}=0.1$. For the LV and NN models, $p_{J}$ is Gaussian; for the SIS model, $p_{J}$ is uniform. Solid lines are solutions of Eq. (7) using the population dynamics algorithm with $5\times 10^{4}$ paths. Symbols denote numerical simulations of Eq. (1) for an ensemble of $10$ independent random networks generated from the configuration model with $N=4000$ nodes. Vertical bars are the standard deviation of the macroscopic observables.
From now on, we set $\sigma=0$ in Eq. (7) and study the dynamics in the absence of external noise. Below we present results for Poisson and geometric indegrees, where the indegree distribution $p_{{\rm in},k}$ is given by
@@
respectively.
Equation (7) is formally similar to other distributional equations appearing in the study of sparse disordered systems Mézard and Parisi (2001, 2003); Kühn (2008); Metz et al. (2019); Abou-Chacra et al. (1973); Mézard and Montanari (2009). Therefore, we can numerically solve this equation using the population dynamics algorithm Mézard and Parisi (2001); Kühn (2008); Metz et al. (2019). In the standard version of this algorithm Mézard and Parisi (2001); Metz et al. (2019), a probability density is parametrized by a population of stochastic variables. Here, we generalize the algorithm to calculate $\mathcal{P}[x]$ by introducing a population of dynamical trajectories. At each iteration step, a single path is chosen randomly from the population and updated according to the differential equation imposed by the Dirac-$\delta_{F}$ in Eq. (7). After sufficient iterations, the population of paths converges to a stationary distribution, providing a numerical solution for $\mathcal{P}[x]$. A detailed account of the algorithm is in Metz .
In Fig. 1, we compare the solutions of Eq. (7) with numerical simulations of the original coupled dynamics, Eq. (1), for finite $N$. The panels showcase the time-evolution of the mean $m(t)$ and the standard deviation $\sqrt{q(t)-m^{2}(t)}$ of the microscopic variables for different network topologies and three important models of complex systems (see table 1): the NN model of Sompolinsky et al. (1988), the LV model, and the SIS model of epidemic spreading. In all cases, the agreement between our theoretical results for $N\rightarrow\infty$ and numerical simulations for large $N$ is excellent, confirming the exactness of Eq. (7).
An important question is whether Eq. (7) recovers the analytic results of fully-connected models as $c\rightarrow\infty$ Sompolinsky et al. (1988); Crisanti and Sompolinsky (2018); Galla (2018). We show in Metz that this is generally not the case. In the high-connectivity limit $c\rightarrow\infty$, degree fluctuations remain significant, and $\mathcal{P}[x]$ depends on the full distribution $\nu_{\rm in}(\kappa)$ of rescaled indegrees $\kappa_{i}=K_{i}/c$ ($i=1,\dots,N$). The well-known effective dynamics on fully-connected networks is only recovered when $\nu_{\rm in}(\kappa)=\delta(\kappa-1)$. This breakdown in the universality of fully-connected models due to degree fluctuations was anticipated in Metz and Silva (2020); Metz and Peron (2022); Silva and Metz (2022); Ferreira and Metz (2023).
@@
Now we derive the phase diagram of the NN model on sparse directed networks for $\sigma=0$. In this context, $x_{i}(t)\in\mathbb{R}$ represents the synaptic current at neuron $i$, $f(x)=x$, and $g(x,x^{\prime})=\tanh{\left(x^{\prime}\right)}$. Let us order the complex eigenvalues $\{\lambda_{\alpha}\}_{\alpha=1,\dots,N}$ of $\boldsymbol{A}$ according to their real parts as ${\rm Re}\lambda_{1}>{\rm Re}\lambda_{2}>\dots>{\rm Re}\lambda_{N}$. A linear stability analysis of Eq. (1) shows that $\boldsymbol{x}=0$ is stable if ${\rm Re}\lambda_{1}<1$. Based on previous analytic results for the spectra of directed networks for $N\rightarrow\infty$ Neri and Metz (2016, 2020); Metz and Neri (2021), we find that the trivial fixed-point is stable provided $c<c_{\rm stab}$, where
@@
The two distinct regimes in Eq. (10) result from the gap-gapless transition in the spectrum of $\boldsymbol{A}$ Neri and Metz (2016). For $c>1+\sigma_{J}^{2}/\mu_{J}^{2}$, the spectral gap $|\lambda_{1}-\lambda_{2}|$ remains finite as $N\rightarrow\infty$, while it vanishes for $c\leq 1+\sigma_{J}^{2}/\mu_{J}^{2}$. Equation (10) holds for $\mu_{J}>0$ and $c>1$. The later condition ensures that directed random networks contain a giant strongly connected component Timár et al. (2017).

### v2

@@ Introduction.
Complex dynamical systems are modeled by $N$ degrees of freedom $x_{i}(t)$ ($i=1,\dots,N$) that evolve in time according to the differential equation
$\dot{x}_{i}(t)=-f(x_{i})+\sum_{j=1}^{N}A_{ij}g(x_{i},x_{j}),$ (1)
where $\{A_{ij}\}_{i,j=1,\dots,N}$ defines the interaction network. The function $f(x)$ governs the dynamics in the absence of interactions, and the kernel $g(x,x^{\prime})$ shapes the pairwise couplings. Equation (1) models the nonequilibrium dynamics of neural networks Sompolinsky et al. (1988); Kadmon and Sompolinsky (2015); Martí et al. (2018); Crisanti and Sompolinsky (2018); Martorell et al. (2024) and ecosystems Opper and Diederich (1992); Bunin (2017); Altieri et al. (2021), epidemic spreading Mata and Ferreira (2013); Mieghem (2012); Goltsev et al. (2012); Pastor-Satorras et al. (2015), synchronization phenomena Stiller and Radons (1998); Rodrigues et al. (2016); Prüser et al. (2024), opinion dynamics Baumann et al. (2020, 2021), and multivariate Ornstein-Uhlenbeck processes Godrèche and Luck (2018); Crisanti and Paoluzzi (2023); Ferraro et al. (2024). Table 1 specifies $f(x)$ and $g(x,x^{\prime})$ for paradigmatic models of complex behavior.
@@
Ornstein-Uhlenbeck process Godrèche and Luck (2018) $x$ $x^{\prime}$
@@
Table 1: Explicit form of $f(x)$ and $g(x,x^{\prime})$ in complex systems modeled by Eq. (1). Equation (1) for the SIS model is derived through the quenched mean-field approximation Pastor-Satorras et al. (2015); Kiss et al. (2017); Silva et al. (2020).
@@
The foremost problem in the study of complex systems is how to reduce the dynamics of many interacting elements to the dynamics of a few variables Thibeault et al. (2024). Dynamical mean-field theory (DMFT) Martin et al. (1973); Coolen (2001); Hertz et al. (2016) is a powerful method to tackle this problem in the limit $N\rightarrow\infty$, yielding a solution in terms of the path-probability for the effective dynamics of a single degree of freedom. The application of DMFT to models described by Eq. (1) has been attracting an enormous interest Kadmon and Sompolinsky (2015); Martí et al. (2018); Crisanti and Sompolinsky (2018); Roy et al. (2019); Martorell et al. (2024); Bunin (2017); Galla (2018); Schuecker et al. (2018); Mastrogiuseppe and Ostojic (2018); Altieri et al. (2021); Prüser et al. (2024); Crisanti and Paoluzzi (2023); Poley et al. (2023); Ferraro et al. (2024); Aguirre-López (2024); Park et al. (2024); Pham and Kaneko (2024), especially in the context of neural networks and ecosystems. Phase diagrams of these models reveal a rich phenomenology, including different types of phase transitions Bunin (2017); Mastrogiuseppe and Ostojic (2018); Martorell et al. (2024), chaotic behavior Sompolinsky et al. (1988); Martí et al. (2018), and coexistence of multiple attractors Bunin (2017); Altieri et al. (2021); Ros et al. (2023). Despite this substantial theoretical progress, DMFT is currently limited to dense networks, where each dynamical variable is essentially coupled to all others by means of Gaussian interaction strengths. However, the interactions in real-world complex systems are known to be sparse and heterogeneous Newman (2010); Guimarães (2020). Sparseness indicates that each element of the system interacts on average with a finite number of others, while heterogeneity refers to fluctuations in the local topology of the interaction network. How to integrate these more realistic features in the formalism of DMFT for systems modeled by Eq. (1) remains an unresolved challenge, and even basic questions, such as deriving the phase diagram of sparse complex systems, are still out of reach.
@@
On the other hand, the dynamics of complex systems on sparse networks has been extensively studied in the case of Ising spins Hatchett et al. (2004); Neri and Bollé (2009); Mimura and Coolen (2009); Roudi and Hertz (2011); Aurell and Mahmoudi (2012); Zhang (2012); Del Ferraro and Aurell (2015); Aurell et al. (2017); Vázquez et al. (2017); Torrisi et al. (2021, 2022); Behrens et al. (2023, 2024). In this context, both the cavity method Neri and Bollé (2009) and DMFT Hatchett et al. (2004); Mimura and Coolen (2009) provide an analytic solution in terms of the path-probability for the effective dynamics of a single dynamical variable. However, the presence of bidirected edges induces a temporal feedback, which, in the particular case of sparse systems, leads to an exponential growth of the path-probability dimension Hatchett et al. (2004); Mimura and Coolen (2009); Behrens et al. (2023), rendering numerical computations unfeasible. Approximation schemes, such as the one-time approximation Neri and Bollé (2009), make these computations possible Zhang (2012); Aurell and Mahmoudi (2012); Torrisi et al. (2022). When the interactions are directed or unidirectional, there is no temporal feedback and the path-probability equation can be efficiently solved Derrida et al. (1987); Hatchett et al. (2004); Neri and Bollé (2009).
@@
Inspired by these results for Ising spins, in this Letter we solve the dynamics of models governed by Eq. (1) on sparse directed networks with an heterogeneous topology. The solution represents a foundational step in the study of complex systems, as it ultimately incorporates the sparse and heterogeneous structure of complex networks into the formalism of DMFT. Directed networks are interesting because they model the non-reciprocal interactions in real-world complex systems Asllani et al. (2018), including the human cortex Peng et al. (2024), food-webs Dunne et al. (2002); Bascompte (2009); Guimarães (2020), gene regulatory networks Shen-Orr et al. (2002); Lee et al. (2002), online social networks Kwak et al. (2010); Schweimer et al. (2022), and the World Wide Web Pastor-Satorras and Vespignani (2004). The formalism presented here thus opens the possibility to analytically investigate how realistic interactions impact the dynamics of complex systems.
@@
By generalizing DMFT to sparse systems, we obtain an exact equation for the path-probability describing the effective dynamics of a single variable for $N\rightarrow\infty$, and we solve this equation for different models in table 1 by using the population dynamics algorithm Mézard and Parisi (2001, 2003); Kühn (2008); Metz et al. (2019); Abou-Chacra et al. (1973); Mézard and Montanari (2009). The excellent agreement between our theoretical results and numerical simulations of finite systems confirms that our solution applies to various nonlinear models interacting through different network topologies, including networks with power-law degree distributions Newman (2010).
@@
As an application, we determine the phase diagram of the neural network model by Sompolinsky et al Sompolinsky et al. (1988) in the sparse regime. We show that the phase diagram displays trivial and nontrivial fixed-point phases, a chaotic phase with zero mean activity, and a chaotic phase with nonzero mean activity Mastrogiuseppe and Ostojic (2018). By calculating certain macroscopic observables, we determine the transition lines as functions of the mean degree and the variance of the coupling strengths, showing their consistency with the universal critical lines derived from random matrix theory Neri and Metz (2016, 2020); Metz and Neri (2021). In particular, we provide numerical evidence that the transition between the chaotic phases coincides with the vanishing of the gap between the leading and the subleading eigenvalue of the interaction matrix.
@@
Let $A_{ij}=C_{ij}J_{ij}$ be the elements of the $N\times N$ interaction matrix $\boldsymbol{A}$. The binary variables $C_{ij}\in\{0,1\}$ determine the network topology, while $J_{ij}\in\mathbb{R}$ controls the interaction strengths. If $C_{ij}=1$, there is a directed edge $j\rightarrow i$ pointing from node $j$ to $i$, while $C_{ij}=0$ otherwise. The indegree $K_{i}=\sum_{j=1}^{N}C_{ij}$ and the outdegree $L_{i}=\sum_{j=1}^{N}C_{ji}$ count the number of links entering and leaving node $i$ Newman (2010), respectively. The random variables $\{C_{ij}\}_{i\neq j}$ follow the distribution
@@
where $\mathcal{N}$ is the normalization constant and $C_{ii}=0$. The degrees $\{K_{i},L_{i}\}_{i=1,\dots,N}$ are independent and identically distributed random variables drawn from $p_{k,\ell}=p_{{\rm in},k}\,p_{{\rm out},\ell}$, where $p_{{\rm in},k}$ and $p_{{\rm out},\ell}$ are the indegree and the outdegree distribution Newman (2010); Metz et al. (2019), respectively. The parameter $c$ is the average degree
@@
Equation (2) defines a network ensemble where directed links are randomly placed between pairs of nodes with probability $c/N$, subject to the prescribed degree sequences generated from $p_{k,\ell}$. In the limit $N\rightarrow\infty$, network samples generated from Eq. (2) are similar to those produced by the configuration model Newman et al. (2001); Fosdick et al. (2018). The coupling strengths $\{J_{ij}\}_{i,j=1,\dots,N}$ are independent and identically distributed random variables drawn from a distribution $p_{J}$ with mean $\mu_{J}$ and variance $\sigma_{J}^{2}$. The distributions $p_{k,\ell}$ and $p_{J}$ fully specify the network ensemble, allowing for a systematic investigation of how network heterogeneities impact the dynamics of complex systems.
@@
of the probability density $\mathcal{P}[\boldsymbol{x}]$ of observing a dynamical path of states $\boldsymbol{x}(t)=(x_{1}(t),\dots,x_{N}(t))$ in a fixed time interval. The correlation functions of $\{x_{i}(t)\}_{i=1,\dots,N}$ follow from the derivatives of $\mathcal{Z}\left[\boldsymbol{\psi}\right]$ with respect to the external sources $\boldsymbol{\psi}(t)=(\psi_{1}(t),\dots,\psi_{N}(t))$. The $n$th-moment of the local variable $x_{i}(t)$,
@@
Clearly, $\mathcal{Z}\left[\boldsymbol{\psi}\right]$ fulfills $\mathcal{Z}\left[0\right]=1$.
In  [^1] See the Supplemental Material for all derivations of DMFT and a detailed account of the population dynamics algorithm., we calculate the average of $\mathcal{Z}\left[\boldsymbol{\psi}\right]$ over the network ensemble defined by Eq. (2) for finite $c$, recasting the problem in terms of the solution of a saddle-point integral. More importantly, we give a clear physical interpretation of the order-parameters, simplifying the saddle-point equations and obtaining a feasible solution for the dynamics in the sparse regime. In the limit $N\rightarrow\infty$, the microscopic dynamical variables decouple, and the path-probability $\mathcal{P}[x]$ for the effective dynamics of a single variable $x(t)$ is determined from
@@
×$\times\delta_{F}\Bigg{[}\dot{x}(t)+f(x(t))-\sum\limits_{j=1}^{k}J_{j}\,g(x(t),x_{j}(t))\Bigg{]},$ (7)
where $\delta_{F}$ is the functional Dirac-$\delta$. The macroscopic observables are computed from $m(t)=\langle x(t)\rangle_{*}$ and $q(t)=\langle x^{2}(t)\rangle_{*}$, where
@@
is the average over the effective dynamics governed by $\mathcal{P}[x]$. The self-consistent Eq. (7) is the exact solution of a broad class of models (see table 1) on sparse directed networks with a local treelike structure Bordenave and Lelarge (2010) and arbitrary distributions $p_{J}$ and $p_{k,\ell}=p_{{\rm in},k}p_{{\rm out},\ell}$. In Note (1), we address the more general case of dynamical models with Gaussian additive noise on networks with correlated indegrees and outdegrees. The solution of Eq. (7) determines the time-evolution of the full probability distribution of the microscopic variables in the limit $N\rightarrow\infty$.
Figure 1: Dynamics of the mean $m(t)$ and the standard deviation $\sqrt{q(t)-m^{2}(t)}$ for the SIS model, LV model, and NN model (see table 1). Results are shown for directed random networks with average degree $c=5$ and two indegree distributions (Eq. (9)): Poisson (left column) and geometric (right column). The distribution $p_{J}$ of the coupling strengths has mean $\mu_{J}$ and standard deviation $\sigma_{J}=0.1$. For the LV and NN models, $p_{J}$ is Gaussian; for the SIS model, $p_{J}$ is uniform. Solid lines are solutions of Eq. (7) using the population dynamics algorithm with $5\times 10^{4}$ paths and initial condition $x_{i}(0)=10^{-3}$. Symbols denote numerical simulations of Eq. (1) for an ensemble of $10$ random networks generated from the configuration model with $N=4000$ nodes. Vertical bars are the standard deviation of the macroscopic observables.
Equation (7) is formally similar to other distributional equations appearing in the study of sparse disordered systems Mézard and Parisi (2001, 2003); Kühn (2008); Metz et al. (2019); Abou-Chacra et al. (1973); Mézard and Montanari (2009). Therefore, we can numerically solve this equation using the population dynamics algorithm Mézard and Parisi (2001); Kühn (2008); Metz et al. (2019). In the standard version of this algorithm Mézard and Parisi (2001); Metz et al. (2019), a probability density is parametrized by a population of stochastic variables. Here, we generalize the algorithm to calculate $\mathcal{P}[x]$ by introducing a population of dynamical trajectories. At each iteration step, a single path is chosen randomly from the population and updated according to the differential equation imposed by the Dirac-$\delta_{F}$ in Eq. (7). After sufficient iterations, the population of paths converges to a stationary distribution, providing a numerical solution for $\mathcal{P}[x]$. A detailed account of the algorithm is in Note (1).
In Fig. 1, we compare the solutions of Eq. (7) with numerical simulations of the original coupled dynamics, Eq. (1), for finite $N$. The panels showcase the time-evolution of the mean $m(t)$ and the standard deviation $\sqrt{q(t)-m^{2}(t)}$ of the microscopic variables for different network topologies and three models of table 1: the NN model of Sompolinsky et al. (1988), the LV model, and the SIS model of epidemic spreading. The results in Fig. 1 are for Poisson and geometric indegrees, where $p_{{\rm in},k}$ is given by
@@
respectively. In Note (1), we compare the solutions of Eq. (7) with numerical simulations for two additional cases: the NN model on networks with power-law indegree distributions and the SIS model at the epidemic threshold. In all cases, the agreement between our theoretical results for $N\rightarrow\infty$ and numerical simulations for large $N$ is excellent, confirming the exactness of Eq. (7).
An important question is whether Eq. (7) recovers the analytic results of fully-connected models as $c\rightarrow\infty$ Sompolinsky et al. (1988); Crisanti and Sompolinsky (2018); Galla (2018); Mastrogiuseppe and Ostojic (2018). By taking the limit $c\rightarrow\infty$ after the thermodynamic limit $N\rightarrow\infty$, we are effectively considering the scaling regime where $c\propto N^{a}$, with $10<a<$ 1 Metz and Silva (2020); Metz and Peron (2022); Silva and Metz (2022); Ferreira and Metz (2023). We show in Note (1) that, in the limit $c\rightarrow\infty$, degree fluctuations remain significant, and $\mathcal{P}[x]$ depends on the full distribution $\nu_{\rm in}(\kappa)$ of rescaled indegrees $\kappa_{i}=K_{i}/c$ ($i=1,\dots,N$). The well-known effective dynamics on fully-connected networks is only recovered when $\nu_{\rm in}(\kappa)=\delta(\kappa-1)$. This breakdown in the universality of fully-connected models due to degree fluctuations was anticipated in Metz and Silva (2020); Metz and Peron (2022); Silva and Metz (2022); Ferreira and Metz (2023).
@@
To demonstrate the strength of Eq. (7) and its concrete applications, we derive the phase diagram of the NN model on sparse directed networks. In this context, $x_{i}(t)\in\mathbb{R}$ represents the synaptic current at neuron $i$, $f(x)=x$, and $g(x,x^{\prime})=\tanh{\left(x^{\prime}\right)}$. Let us order the complex eigenvalues $\{\lambda_{\alpha}\}_{\alpha=1,\dots,N}$ of $\boldsymbol{A}$ according to their real parts as ${\rm Re}\lambda_{1}\geq{\rm Re}\lambda_{2}\geq\dots\geq{\rm Re}\lambda_{N}$. A linear stability analysis of Eq. (1) shows that $\boldsymbol{x}=0$ is stable if ${\rm Re}\lambda_{1}<1$. Based on analytic results for the spectra of directed networks for $N\rightarrow\infty$ Neri and Metz (2016, 2020); Metz and Neri (2021), we find that the trivial fixed-point is stable provided $c<c_{\rm stab}$, where
@@
The two distinct regimes in Eq. (10) result from the gap-gapless transition in the spectrum of $\boldsymbol{A}$ Neri and Metz (2016). For $c>1+\sigma_{J}^{2}/\mu_{J}^{2}$, the spectral gap $|\lambda_{1}-\lambda_{2}|$ remains finite as $N\rightarrow\infty$, while it vanishes for $c\leq 1+\sigma_{J}^{2}/\mu_{J}^{2}$. Equation (10) holds for $\mu_{J}>0$ and $c>1$. The condition $c>1$ ensures that directed networks with degree distribution $p_{k,\ell}=p_{{\rm in},k}p_{{\rm out},\ell}$ contain a giant strongly connected component Timár et al. (2017); Kryven (2016), implying the existence of a continuous part in the eigenvalue distribution of $\boldsymbol{A}$ Neri and Metz (2020).